OpenStudy (anonymous):

Write the equation of the line that passes through (-4,12) and (7,23) A. Y=x+16 B. y=3x+18 C. Y=1/6x+16 D. Y=1/3x+18

OpenStudy (anonymous):

Please help I'm desperate

OpenStudy (kropot72):

The equation is of the form y = mx + c where m is the slope. You can find the slope from \[m=\frac{y _{1}-y _{2}}{x _{1}-x _{2}}\] Can you do that?

OpenStudy (anonymous):

A

OpenStudy (kropot72):

@micahcarr76 You choice is correct.

OpenStudy (anonymous):

The point-slope form of a line is \(y-y_1=m(x-x_1)\). You can find the slope \(m\) using:$$m=\frac{y_2-y_1}{x_2-x_1}=\frac{23-12}{7--4}=\frac{11}{11}=1$$Therefore we can simplify the above point-slope form to \(y-y_1=1(x-x_1)=y-y_1=x-x_1\). We can pick either point to plug into our template for the point-slope form so I'll do both. Let's try \((-4,12)\), which yields:$$y-12=x--4=x+4\\y=x+4+12=x+16$$ To prove that it also works using the other point, let's try \((7,23)\):$$y-23=x-7\\y=x-7+23=x+16$$ So our answer is conclusively (A).

OpenStudy (anonymous):

Already had the answer and was about to close the question when u guys responded